# A Survey on Failure Analysis and Fault Injection in AI Systems

### v1

@@ 7.1. Failure Analysis in AI Toolkit
Synchronization Fault is frequently encountered due to the concurrent nature as GPU programs commonly operate using multiple threads. In contrast to CPUs that commonly utilize lock mechanisms for data synchronization, GPUs predominantly rely on barriers as their synchronization mechanism. In particular, a barrier is represented as a barrier function __syncthreads() in CUDA kernel functions (NVIDIA, 2024a). There are primarily three main causes for synchronization faults: data race, barrier divergence, and redundant barrier function  (Collingbourne et al., 2013; Wu et al., 2019). Data race primarily occurs due to the inability to determine the order of ”read&write” and ”write&write” actions among multiple threads (Islam and Muzahid, 2018; Wu et al., 2019). Barrier Divergence occurs when certain threads within the same block fail to reach a barrier due to variations in their execution flow. One common scenario is when a barrier (e.g., __syncthreads()) is positioned inside an if code block, leaving only a subset of threads can reach the barrier. Redundant Barrier Function is typically caused by unnecessary synchronization operations, which can result in inefficiencies in both speed and memory utilization of GPU programs.
@@
Memory Safety Fault is frequently observed in low-level programming languages which provide direct memory access and management, particularly those specifically tailored for GPU programming(e.g., CUDA). Memory safety is typically ensured by the restriction that memory allocations can only be accessed between their intended bounds and during their lifetime (Song et al., 2019; Ziad et al., 2023). The primary causes of memory safety faults can be categorized into three factors: out-of-bounds access, temporal safety, and failed free operation (Ziad et al., 2023). In GPU programming, out-of-bounds access encompasses accessing buffers beyond their boundaries in global memory or shared memory. Temporal safety faults primarily arise when attempting to access GPU memory that has already been freed or accessing GPU memory that has not been properly allocated or initialized. Additionally, use-after-scope fault  (Ziad et al., 2023) can occur in local memory of GPU. Failed free operation includes double free and invalid free. Double free occurs when attempting to free memory that has already been freed, while invalid free refers to freeing memory that was not dynamically allocated.
@@
Dependency Fault primarily occurs when there is a mismatch between the AI toolkit and the higher-level AI framework or AI application. Based on the number of components involved, dependency fault can be categorized into two types (Huang et al., 2023): intra- and inter-dependency fault. Intra-dependency faults can occur due to incorrect versioning or unsuccessful installation of a toolkit. Unsuccessful installation can also be categorized into two types: missing installation of required libraries and incorrect path configuration. Inter-dependency faults in AI toolkit can occur can arise from a mismatch of software or hardware. An example of software mismatch is the mismatch between CUDA version and PyTorch version, which may occur a ”driver too old” fault when running PyTorch (PyTorch, 2024a). Hardware mismatch can occur when the hardware lacks specific features required by the AI framework. For example, TensorFlow 1.6 utilizes the AVX feature of CPUs. However, if the CPU does not support AVX, it can lead to a dependency fault.
@@
Communication Fault primarily occurs within the communication mechanism of distributed AI training. As two mostly common communication mechanisms in distributed AI, the faults occurring in NCCL (NVIDIA, 2024c) and NVLink can significantly impact the workload of distributed AI (Gao et al., 2023b; Hu et al., 2024), e.g., the training of LLM. The nccl fault is possibly due to a network fault or a remote process exiting prematurely (Gao et al., 2023b). And the NVLink fault is mainly caused by the hardware failures in GPU. Hu et al. (Hu et al., 2024) observes that training 7B models in Kalos tend to result in GPU overheating, which can cause NVLink fault. This phenomenon increases with the optimization of communication costs because it leads to more exceptionally low GPU idle rates. Another commonly used communication tool is MPI (mpich, 2024). AI systems that utilize MPI also faces the faults associated with MPI itself. This kind of faults is usually due to either a fault of network connection to peer MPI process, or possibly an internal fault of the MPI daemon itself (Jeon et al., 2019).
@@ 7.2. Fault Injection in AI Toolkits
CUDA. As a parallel computing platform and application programming interface (API) developed by NVIDIA, CUDA offers AI systems high-performance and highly parallel capabilities. Recently, there are two main areas of work in fault injection at the CUDA level.
@@
  • FI in CUDA Compilers. Faults in the CUDA compiler can either lead to compile-time faults or emit executable codes that may lead to runtime faults. CUDAsmith (Jiang et al., 2020) proposes a FI framework for CUDA compilers, which can be used to test several versions of NVCC and Clang compilers for CUDA with different optimization levels. CUDAsmith continuously generates CUDA programs to detect potential bugs in the CUDA compiler and utilizes equivalence modulo inputs (EMI) testing techniques to solve the test oracle problem.
@@
Collective Communication. Collective communication is defined as communication that involves a group of processes, which plays a significant role in distributed AI scenarios. The intricate communication among different nodes poses significant challenges to the reliability of collective communication. FastFIT (Feng et al., 2015) is a fault injection tool specifically designed for MPI. It injects faults randomly into the input parameters of collective interface. In particular, FastFIT manifests the fault by one bit flip in one of the input parameters, which typically include the send/receive buffer address, data elements, data type, communication destination, and communicator.
@@
Failure Simulee CUDAsmith CLsmith FastFIT Covered
Data Race True
@@
Out-of-Bounds Access True
@@
Inter-dependency Fault True
@@ 7.3. Gap between Failure Analysis and Fault Injection in AI Tookit
Incomplete coverage of fault types. Table 12 highlights that certain types of faults, such as ”Temporal Safety Fault”, ”Failed Free Operation”, ”Intra-dependency Fault”, ”NCCL Fault”, and ”NVLink Fault”, are not covered by the existing FI tools listed. This gap indicates that the current FI techniques may not comprehensively address the diverse range of faults that can occur in AI toolkits, potentially leaving blind spots in testing and resilience evaluation.
Lack of FI capabilities for distributed and parallel computing. While tools like FastFIT can inject faults related to MPI and inter-dependency faults, there is a lack of FI capabilities specifically targeting faults that can arise in distributed and parallel computing environments. AI toolkits increasingly leverage distributed and parallel computing for efficient model training and inference, making it crucial to have FI techniques that can simulate and analyze faults in these scenarios, such as ”NCCL Fault” and ”NVLink Fault”.
@@ 8.1. Failure Analysis in AI Platform
In this section, we primarily introduce FA about Spark (Zaharia et al., 2010), Ray (Moritz et al., 2018) and Platform-X in Microsoft (Gao et al., 2023b), which are three representative AI platforms. Due to limited FA work in platform layer, we supplement several fault types based on the merged pull requests (PRs) that are responsible to fix bugs on GitHub (Ray-project, 2024). Notably, not all pull requests in this category are exclusively for bug fixes. Some may focus on introducing new features or updating documentation. To specifically identify bug-fixing pull requests, we employ keyword searches in the tags and titles, leveraging established bug-related terms such as fix, defect, fault, bug, issue, mistake, correct, fault, and flaw, aligning with prior research (Islam et al., 2019; Shen et al., 2021). Table 13 shows the detailed failures of AI platforms.
@@
Memory Leak Unreleased memory causing system slowdown. (Ray, 2021a; Spark, 2021)
@@
Resource Overload Excessive resource usage leading to system overload. (Gao et al., 2023b; Zhang et al., 2020b)
@@
Platform Resource Faults include resource contention and resource overload problem. Resource contention occurs when multiple components or workloads compete for shared resources, leading to performance bottlenecks (Gao et al., 2023b; Shah et al., 2021; Lu et al., 2017). Sarah et al. (Shah et al., 2021) and Lu et al. (Lu et al., 2017) analyze the performance impact caused by interference between Spark applications from the perspective of mutual interference and propose a technology that can quickly diagnoses the root cause of interference. Resource overload, on the other hand, refers to situations where excessive resource usage causes system overload and performance degradation (Gao et al., 2023b; Zhang et al., 2020b).
@@ 8.2. Fault Injection in AI Platform
As a distributed computing architecture, the communication between nodes and the influence between node states are commom faults in AI platforms, such as node partition caused by network fault and node recovery bugs caused by node crashes. Therefore, there are currently some related works that inject faults into these issues to discover corresponding bugs. ChaosBlade (Chaosblade, 2024) can introduce resource hog faults into target system to test its resilience. Chen et al. propose a consistency-guided fault injection technique called CoFI to systematically injects network partitions to effectively expose partition bugs in distributed systems (Chen et al., 2020a). Gao et al. propose CrashFuzz, a fault injection testing approach that can effectively test crash recovery behaviors and reveal crash recovery bugs in distributed systems (Gao et al., 2023a).
@@ 8.3. Gap between Failure Analysis and Fault Injection in AI Platform
Limited Coverage. Not all types of failures are covered by the listed FI tools. For instance, failures due to misconfiguration, wrong access control, memory leaks, tool/library faults, and GPU resource contention are not being simulated by any of the tools. This indicates a significant gap in the current FI capabilities and highlights areas where further research are needed.
Lack of Specific Design. The current fault injection tools appear to be designed primarily for traditional cloud platforms, such as Kubernetes, with less consideration given to the unique characteristics and requirements of AI platforms. For instance, none of the listed tools can simulate failures related to GPU resource contention, which is a critical aspect of AI platforms due to their heavy reliance on GPU resources for computation-intensive tasks. This lack of specific design for AI platforms introduces a substantial gap in the ability to accurately simulate and study the full range of potential failures in these systems.
Failure ChaosBlade CoFI CrashFuzz TRANSMUT-SPARK DepFuzz Coverd
Concurrency Fault True
API Incompatibility True
Misconfiguration False
Wrong Access Control False
Exception Fault True
Memory Leak False
Tool/Library Fault False
Misoperation True
Resource Contention (excl. GPU) True
Resource Contention (GPU) False
Resource Overload False
@@ 9. Failure Analysis and Fault Injection for AI Infrastructure
AI infrastructure layer serves as the foundational layer in an AI system architecture, providing the underlying physical and virtualized resources necessary for the deployment and operation of AI applications and services. This layer is responsible for managing and orchestrating the computing, storage, and networking resources required by the AI platform layer and other higher-level layers. Potential failures in this layer can lead to service unavailability, system deployment failures, model training failures, etc. This section delves into the FA and FI in AI infrastructure layer.
@@ 9.1.1. Hardware Accelerators
We analyze the failure of GPU, FPGA and TPU, which are three representative hardware accelerators.
@@
GPU. GPU has become the most commonly used underlying hardware in AI and HPC systems. However, according to existing research, the frequency of failures caused by GPU is still high (Taherin et al., 2021). Research conducted on the Titan supercomputer explores different aspects of GPU failures. This includes the examination of GPU faults in a broader context (Tiwari et al., 2015a), the analysis of specific GPU software faults (Nie et al., 2016), the characterization of GPU failures concerning temperature and power (Nie et al., 2017), and the investigation of spatial characteristics associated with failures (Tiwari et al., 2015b). Ostrouchov et al.  (Ostrouchov et al., 2020) find that GPU reliability is dependent on heat dissipation to an extent that strongly correlates with detailed nuances of the cooling architecture and job scheduling. Nie et al. (Nie et al., 2017) analyze the relationship between single bit faults occurrence and temperature on the Titan supercomputer, and propose a machine learning based technique for GPU soft-fault prediction. A study about another supercomputer, Blue Water, analyzes GPU failures among other hardware failures (Di Martino et al., 2014). The study reveals that GPUs rank among the top three most prone to failures, and notably, GPU memory exhibits greater sensitivity to uncorrectable faults compared to main memory.
@@
Given the distinctions in workload between HPC and AI systems, the following discussion delves into GPU failure analysis specifically tailored to AI systems. Zhang et al. (Zhang et al., 2020b) present the first comprehensive empirical study on program failures of deep learning jobs and found that the GPU ”Out of Memory” fault accounts for 65.0% of the failures in the deep learning specific dimension. Since in a large-scale deep learning cluster, GPU failures are inevitable and they cause severe consequences, Liu et al. (Liu et al., 2023c) propose prediction models of GPU failures under large-scale production deep learning workloads. The prediction model takes into account static parameters such as GPU type, as well as dynamic parameters such as GPU temperature and power consumption, and integrates parallel and cascading architectures to make good predictions of GPU failures.
@@
FPGA. FPGA is a digital technology designed to be configured by a customer or a designer after manufacturing, hence the term ”field-programmable”. As a neural network accelerator, FPGA is the subject of various studies related to reliability. Radu et al. (Radu, 2014) propose a new probabilistic method, the Component Failure Analysis (CFA), that uses FPGA specific techniques and algorithms for analyzing SEUs in implemented FPGA designs. McNelles et al. (McNelles and Lu, 2016) use Dynamic Flowgraph Methodology (DFM) to model FPGA, showing the potential advantage of DFM for modeling FPGA-based systems compared with static methods and simulation.
@@
Examining an FPGA from diverse perspectives leads to varied insights and advantages. Conmy et al. (Conmy and Bate, 2010) employ a semi-automated FPTC analysis technique, customized for specific fault types identified on an FPGA, to thoroughly examine individual faults within electronic components. These components support a modularized design embedded on the FPGA. The study demonstrates how the analysis of these individual faults can be seamlessly integrated with crosscutting safety analysis, thereby reinforcing and validating the necessary safety properties. Xu et al. (Xu et al., 2021) take the entire FPGA-based neural network accelerator, including the control unit and DMA modules into consideration. The experiments on four typical neural networks showed that hardware faults can incur both system exceptions, such as system stall and prediction accuracy loss.
@@
TPU. TPU is initially developed by Google to accelerate machine learning workloads, specifically targeting the training and inference of deep neural networks within the TensorFlow framework. Pablo et al. (Bodmann and Rech, 2024) measure TPU’s atmospheric neutrons reliability at different temperatures, that goes from -40°C to +90°C. They show a decrease in the FIT rate of almost 4× as temperature increases. Rubens et al. (Junior and Rech, 2022) investigat the reliability of TPU executing 2D-3D convolutions and eight CNNs to high-energy, mono-energetic, and thermal neutron. They find that despite the high fault rate, most neutron-induced faults do not change the CNNs detection/classification. Rubens et al. (Junior et al., 2022) investigate the reliability of TPUs to atmospheric neutrons, reporting experimental data equivalent to more than 30 million years of natural irradiation.

### v2

@@ 7.1. Failure Analysis in AI Toolkit
Synchronization Fault is frequently encountered due to the concurrent nature as GPU programs commonly operate using multiple threads. In contrast to CPUs that commonly utilize lock mechanisms for data synchronization, GPUs predominantly rely on barriers as their synchronization mechanism. In particular, a barrier is represented as a barrier function __syncthreads() in CUDA kernel functions (NVIDIA, 2024a). There are primarily three main causes for synchronization faults: data race, barrier divergence, and redundant barrier function  (Collingbourne et al., 2013; Wu et al., 2019). Data race occurs primarily due to the inability to determine the order of “read&write” and “write&write” actions among multiple threads (Islam and Muzahid, 2018; Wu et al., 2019). Barrier divergence occurs when certain threads within the same block fail to reach a barrier due to variations in their execution flow. One common scenario is when a barrier (e.g., __syncthreads()) is positioned inside an if code block, leaving only a subset of threads can reach the barrier. Redundant Barrier Function is typically caused by unnecessary synchronization operations, which can result in inefficiencies in both speed and memory utilization of GPU programs.
@@
Memory Safety Fault is frequently observed in low-level programming languages that provide direct memory access and management, particularly those specifically tailored for GPU programming (e.g., CUDA). Memory safety is typically ensured by the restriction that memory allocations can only be accessed between their intended bounds and during their lifetime (Song et al., 2019; Ziad et al., 2023). The primary causes of memory safety faults can be categorized into three factors: out-of-bounds access, temporal safety, and failed free operation (Ziad et al., 2023). In GPU programming, out-of-bounds access encompasses accessing buffers beyond their boundaries in global memory or shared memory. Temporal safety faults arise primarily when attempting to access GPU memory that has already been freed or accessing GPU memory that has not been properly allocated or initialized. Additionally, a use-after-scope fault  (Ziad et al., 2023) can occur in the local memory of the GPU. Failed free operation includes double free and invalid free. Double free occurs when attempting to free memory that has already been freed, while invalid free refers to freeing memory that was not dynamically allocated.
@@
Dependency Fault occurs primarily when there is a mismatch between the AI toolkit and the higher-level AI framework or AI application. Based on the number of components involved, dependency fault can be categorized into two types (Huang et al., 2023): intra- and inter-dependency fault. Intra-dependency faults can occur due to incorrect versioning or unsuccessful installation of a toolkit. Unsuccessful installation can also be categorized into two types: missing installation of required libraries and incorrect path configuration. Inter-dependency faults in AI toolkit can occur can arise from a mismatch of software or hardware. An example of software mismatch is the mismatch between the CUDA version and the PyTorch version, which may occur a “driver too old” fault when running PyTorch (PyTorch, 2024a). Hardware mismatch can occur when the hardware lacks specific features required by the AI framework. For example, TensorFlow 1.6 utilizes the AVX feature of CPUs. However, if the CPU does not support AVX, it can lead to a dependency fault.
@@
Communication Fault primarily occurs within the communication mechanism of distributed AI training. As two mostly common communication mechanisms in distributed AI, the faults occurring in NCCL (NVIDIA, 2024c) and NVLink can significantly impact the workload of distributed AI (Gao et al., 2023b; Hu et al., 2024), e.g., the training of LLM. The nccl fault is possibly due to a network fault or a remote process exiting prematurely (Gao et al., 2023b). The NVLink fault is mainly caused by the hardware failures in GPU. Hu et al. (Hu et al., 2024) observe that training 7B models in Kalos tends to result in GPU overheating, which can cause NVLink fault. This phenomenon increases with the optimization of communication costs because it leads to more exceptionally low GPU idle rates. Another commonly used communication tool is MPI (mpich, 2024). AI systems that utilize MPI also faces the faults associated with MPI itself. This kind of faults is usually due to either a fault of network connection to peer MPI process, or possibly an internal fault of the MPI daemon itself (Jeon et al., 2019).
@@ 7.2. Fault Injection in AI Toolkits
CUDA. As a parallel computing platform and application programming interface (API) developed by NVIDIA, CUDA offers high-performance and highly parallel capabilities for AI systems. Recently, there are two main areas of work in fault injection at the CUDA level.
@@
  • FI in CUDA Compilers. Faults in the CUDA compiler can either lead to compile-time faults or emit executable codes that may lead to runtime faults. CUDAsmith (Jiang et al., 2020) proposes an FI framework for CUDA compilers, which can be used to test several versions of the NVCC and Clang compilers for CUDA with different optimization levels. CUDAsmith continuously generates CUDA programs to detect potential bugs in the CUDA compiler and utilizes equivalence modulo inputs (EMI) testing techniques to solve the test oracle problem.
@@
Collective Communication. Collective communication is defined as communication that involves a group of processes, which plays a significant role in distributed AI scenarios. Intricate communication among different nodes poses significant challenges to the reliability of collective communication. FastFIT (Feng et al., 2015) is a fault injection tool specifically designed for MPI. It injects faults randomly into the input parameters of the collective interface. In particular, FastFIT manifests the fault by one bit flip in one of the input parameters, which typically include the send/receive buffer address, data elements, data type, communication destination, and communicator.
@@
Group Failure Simulee CUDAsmith CLsmith FastFIT Covered
Synchronization Data Race True
@@
Memory Safety Out-of-Bounds Access True
@@
Communication MPI Fault True
NVLink Fault False
@@ 7.3. Gap between Failure Analysis and Fault Injection in AI Tookit
Incomplete coverage of fault types. Table 13 highlights that certain types of fault, such as “Temporal Safety Fault”, “Failed Free Operation”, “Intra-dependency Fault”, “NCCL Fault”, and “NVLink Fault”, are not covered by the existing FI tools listed. This gap indicates that current FI techniques may not comprehensively address the diverse range of faults that can occur in AI toolkits, potentially leaving blind spots in testing and resilience evaluation.
Lack of FI capabilities for distributed and parallel computing. While tools like FastFIT can inject faults related to MPI and inter-dependency faults, there is a lack of FI capabilities specifically targeting faults that can arise in distributed and parallel computing environments. AI toolkits increasingly leverage distributed and parallel computing for efficient model training and inference, making it crucial to have FI techniques that can simulate and analyze faults in these scenarios, such as “NCCL Fault” and “NVLink Fault”.
@@ 8.1. Failure Analysis in AI Platform
In this section, we primarily introduce FA about Spark (Zaharia et al., 2010), Ray (Moritz et al., 2018) and Platform-X in Microsoft (Gao et al., 2023b), which are three representative AI platforms. Due to limited FA work in the platform layer, we supplement several fault types based on the merged pull requests (PRs) that are responsible to fix bugs on GitHub (Ray-project, 2024). Notably, not all pull requests in this category are exclusively for bug fixes. Some may focus on introducing new features or updating the documentation. To specifically identify bug-fixing pull requests, we employ keyword searches in the tags and titles, leveraging established bug-related terms such as fix, defect, fault, bug, issue, mistake, correct, fault, and flaw, aligning with prior research (Islam et al., 2019; Shen et al., 2021). Table 14 shows the detailed failures of the AI platforms.
@@
Memory Leak Unreleased memory causing system slowdown. (Ray, 2021a; Spark, 2021; Du et al., 2020)
@@
Resource Overload Excessive resource usage leading to system overload. (Gao et al., 2023b; Zhang et al., 2020b; Liu et al., 2022)
@@
Platform Resource Faults include resource contention and resource overload problem. Resource contention occurs when multiple components or workloads compete for shared resources, leading to performance bottlenecks (Gao et al., 2023b; Shah et al., 2021; Lu et al., 2017). Sarah et al. (Shah et al., 2021) and Lu et al. (Lu et al., 2017) analyze the performance impact caused by interference between Spark applications from the perspective of mutual interference and propose a technology that can quickly diagnose the root cause of interference. Resource overload, on the other hand, refers to situations where excessive resource usage causes system overload and performance degradation (Gao et al., 2023b; Zhang et al., 2020b).
@@ 8.2. Fault Injection in AI Platform
As a distributed computing architecture, the communication between nodes and the influence between node states are common faults in AI platforms, such as node partition caused by network fault and node recovery bugs caused by node crashes. Therefore, there are currently some related works that inject faults into these issues to discover corresponding bugs. ChaosBlade (Chaosblade, 2024) can introduce resource hog faults into target system to test its resilience. Chen et al. propose a consistency-guided fault injection technique called CoFI to systematically injects network partitions to effectively expose partition bugs in distributed systems (Chen et al., 2020a). Gao et al. propose CrashFuzz, a fault injection testing approach that can effectively test crash recovery behaviors and reveal crash recovery bugs in distributed systems (Gao et al., 2023a).
Group Failure ChaosBlade CoFI CrashFuzz TRANSMUT DepFuzz Coverd
Code Concurrency Fault True
API Incompatibility True
Exception Fault True
Misconfiguration False
Wrong Access Control False
Memory Leak False
Platform Maintenance Tool/Library Fault False
Misoperation True
Platform Resource Resource Contention (excl. GPU) True
Resource Contention (GPU) False
Resource Overload False
@@ 8.3. Gap between Failure Analysis and Fault Injection in AI Platform
Limited Coverage. Not all types of failure are covered by the listed FI tools. For instance, failures due to misconfiguration, wrong access control, memory leaks, tool/library faults, and GPU resource contention are not being simulated by any of the tools. This indicates a significant gap in the current FI capabilities and highlights areas where further research are needed.
Lack of Specific Design. Current fault injection tools appear to be designed primarily for traditional cloud platforms, such as Kubernetes, with less consideration given to the unique characteristics and requirements of AI platforms. For instance, none of the listed tools can simulate failures related to GPU resource contention, which is a critical aspect of AI platforms due to their heavy reliance on GPU resources for computation-intensive tasks. This lack of a specific design for AI platforms introduces a substantial gap in the ability to accurately simulate and study the full range of potential failures in these systems.
@@ 9. Failure Analysis and Fault Injection for AI Infrastructure
AI infrastructure layer serves as the foundational layer in an AI system architecture, providing the underlying physical and virtualized resources necessary for the deployment and operation of AI applications and services. This layer is responsible for managing and orchestrating the computing, storage, and networking resources required by the AI platform layer and other higher-level layers. Potential failures in this layer can lead to service unavailability, system deployment failures, model training failures, etc. This section delves into FA and FI in the AI infrastructure layer.
@@ 9.1.1. Hardware Accelerators
We analyze the failure of GPU, FPGA, and TPU, which are three representative hardware accelerators.
@@
GPU. GPU has become the most commonly used underlying hardware in AI and HPC systems. However, according to existing research, the frequency of failures caused by GPU is still high (Taherin et al., 2021). Research conducted on the Titan supercomputer explores different aspects of GPU failures. This includes the examination of GPU faults in a broader context (Tiwari et al., 2015a), the analysis of specific GPU software faults (Nie et al., 2016), the characterization of GPU failures concerning temperature and power (Nie et al., 2017), and the investigation of spatial characteristics associated with failures (Tiwari et al., 2015b). Ostrouchov et al.  (Ostrouchov et al., 2020) find that GPU reliability is dependent on heat dissipation to an extent that strongly correlates with detailed nuances of the cooling architecture and job scheduling. Nie et al. (Nie et al., 2017) analyze the relationship between single bit faults occurrence and temperature on the Titan supercomputer, and propose a machine learning-based technique for GPU soft-fault prediction. A study about another supercomputer, Blue Water, analyzes GPU failures among other hardware failures (Di Martino et al., 2014). The study reveals that GPUs rank among the top three most prone to failures and GPU memory exhibits greater sensitivity to uncorrectable faults compared to main memory.
@@
Given the distinctions in workload between HPC and AI systems, the following discussion delves into GPU failure analysis specifically tailored to AI systems. Zhang et al. (Zhang et al., 2020b) present the first comprehensive empirical study on program failures of deep learning jobs and found that the GPU “Out of Memory” fault accounts for 65.0% of the failures in the deep learning specific dimension. Since in a large-scale deep learning cluster, GPU failures are inevitable and cause severe consequences, Liu et al. (Liu et al., 2023c) propose prediction models of GPU failures under large-scale production deep learning workloads. The prediction model takes into account static parameters such as GPU type, as well as dynamic parameters such as GPU temperature and power consumption, and integrates parallel and cascading architectures to make good predictions of GPU failures.
@@
FPGA. FPGA is a digital technology designed to be configured by a customer or designer after manufacturing, hence the term “field-programmable”. As a neural network accelerator, FPGA is the subject of various studies related to reliability. Radu et al. (Radu, 2014) propose a new probabilistic method, Component Failure Analysis (CFA), which uses FPGA specific techniques and algorithms for analyzing SEUs in implemented FPGA designs. McNelles et al. (McNelles and Lu, 2016) use Dynamic Flowgraph Methodology (DFM) to model FPGA, showing the potential advantage of DFM for modeling FPGA-based systems compared to static methods and simulation.
@@
Examining an FPGA from diverse perspectives leads to varied insights and advantages. Conmy et al. (Conmy and Bate, 2010) employ a semi-automated FPTC analysis technique, customized for specific fault types identified in an FPGA, to thoroughly examine individual faults within electronic components. These components support a modularized design embedded in the FPGA. The study demonstrates how the analysis of these individual faults can be seamlessly integrated with crosscutting safety analysis, thereby reinforcing and validating the necessary safety properties. Xu et al. (Xu et al., 2021) take the entire FPGA-based neural network accelerator, including the control unit and DMA modules into consideration. The experiments on four typical neural networks showed that hardware faults can incur both system exceptions, such as system stall and prediction accuracy loss.
@@
TPU. TPU is initially developed by Google to accelerate machine learning workloads, specifically targeting deep neural network training and inference within the TensorFlow framework. Pablo et al. (Bodmann and Rech, 2024) measure TPU’s atmospheric neutron reliability at different temperatures that goes from -40°C to +90°C. They show a decrease in the FIT rate of almost 4× as temperature increases. Rubens et al. (Junior and Rech, 2022) investigate the reliability of TPU executing 2D-3D convolutions and eight CNNs to high-energy, mono-energetic, and thermal neutron. They find that despite the high fault rate, most neutron-induced faults do not change the CNNs detection/classification. Rubens et al. (Junior et al., 2022) investigate the reliability of TPUs to atmospheric neutrons, reporting experimental data equivalent to more than 30 million years of natural irradiation.